3) An integer is chosen at random from the first 50bdigits. What is the probability that it is divisible by 6 or 8.
4) An integer is chosen at random from the first 100 positive integers. Find the probability that it is divisible by
i) 8 and 12.
ii) 16 and 20.

Answer by Mathtut(3670)

3) there are only 2 numbers that work 24 and 48 so 2/50=1/25
:
4)
i)only 4 numbers work 24,48,72, and 96 so 4/100=1/25
ii)only 1 number works 80.....so 1/100